Posted by: cheryl searcy (csearcy@usgs.gov), Mon Jan 12, 2026 22:00 (2200 UTC)
--
AVO Meeting notes for January 12, 2026
Great Sitkin: Orange
Shishaldin: Yellow
Ahyi: Yellow
SIC: Matt Haney
CS: David Fee
Duty Scientist: Matt Loewen
AVO/CNMI Duty Remote Sensor: Matt Loewen
AVO/CNMI Duty Seismologist: Cheryl Searcy
Primary alarms: Cheryl Searcy
Secondary alarms: Matt Loewen
AVO radio: Michelle Coombs
DUTY SCIENTIST REPORT– Kristi Wallace
+ GREAT SITKIN (Orange/Watch): No major changes. Cloudy all week. Reston report from Jan 5 confirms lava spreading to the southwest through the southeast with minimal advancement and notes spalling off the dome and lava flow fronts. Weak seismicity occasionally observed.
+ SHISHALDIN (Yellow/Advisory): No major changes. Steaming from the summit was observed in satellite and webcam images when clear. Low-frequency seismic and infrasound activity detected all week when not too windy.
+ KOROVIN (Green/Normal): A discrete seismic event occurred on Jan 6 around 00:43 UTC, followed by elevated tremor. No obvious infrasound but it has been windy. Seismicity (tremor and local VT’s) observed Jan 7, 8, maybe on 10th.
+ SPURR (Green/Normal): There were a few LPs at Spurr last week. They may be glacial.
+ CNMI/AHYI (Yellow/Advisory): Possible Ahyi detection in Wake Island data on Jan 9. Faint submarine plume observed Jan 6, 9 and 10. No activity observed at other CNMI volcanoes.
+ OTHER: Kilauea started event #40 this morning.
NET OPS/FIELDWORK: Max Enders
Color Code:
+ Great Sitkin – All stations working.
+ Shishaldin – All stations working except some SSBA data gaps.
Network, Multi-Station and Array Outages: --
+ Port Heiden (#aniakchak, #veniaminof (north side) - 2025-12-13 07:07:08 UTC Power outage is ONGOING. – Update received from FAA is that a crew is still trying to get to Port Heiden to make repairs, but have not been able to make it yet due to weather.
+ Sand Point (SDPI array) - outage continues, partner-procured internet outage – EarthScope has an open trouble ticket with the ISP, and they are considering a visit to install a Starlink if this is not resolved soon.
+ Kenai Infrasound Array – Outage due to station power issues – Critically low voltage starting ~2025-01-02 21:00 UTC. No field visit planned.
+ Akutan- Network outage due to ISP administrative issue. Out: 2026-01-05 19:50:56 UTC Resolved: 2026-01-07 18:26:01 UTC.
Single Station Issues –
-Wrangell - Tolsona sites continue to have minor data gaps at IRIS. AVO Winston copy complete.
-Iliamna – ILLG is completely out, ILCB very intermittent last week, and ILSW power getting low.
-Peulik – PLBL and R17L intermittent and currently out due to snow/ice at PLBL.
-Hague – D. Fee noted in the SOH channel that low frequency noise worse on the last month.
-Shishaldin – SSBA/AV36 and ISNN/AV38 low power, GNSS powered off at both stations to conserve power for seismic data. -Westdahl – WECS seismic offline to conserve power for GNSS and radios.
-Akutan -LVA out due to low voltage.
-Okmok – OKAK and OKWE no telemetry. –
-Bogoslof – BOGO completely out last week.
-Atka – KOSE/KONE data gaps last week, presumably rime ice at repeater.
-Adak – Infrasound element 6 is not working and element 2 reported as having reversed polarity.
-Tanaga/Gareloi – TASO, GANO and GAEA – Outage early last week likely due to icing/snow on radio antenna at TASE.
-Little Sitkin -Gappy data and outages from LSNW and LSSA probably due to icing on LSSA radio antennas.
Overall**:** None
Tech Team Fieldwork: None
OPEN ANNOUNCEMENTS
+ ADGGS – Scott Crass
• Everyone is at work this week. Israel Brewster is working from home.
• Dain Harmon is sending out a key word survey. Please see his post in the Town Square channel of Mattermost.
+ UAF – David Fee
• The first draft of the CoMeet is out. Please check it over for errors and send corrections/suggestions to David.
• There are several classes that conflict with the Monday morning staff meeting.
+ AVO South – Matt Haney
• Thank you to all who submitted non-net field proposals. There are 13. Scott Crass, David Fee and Matt Haney will be meeting on Wednesday to go over them.
• Next Monday is a holiday. The staff meeting will be moved to Tuesday.
• There have been changs in the VSC leadership staffing effective yesterday:
o Acting VSC director: Jake Lowenstern
o Acting CVO SIC: Seth Moran
o VDAP director: Heather Wright.
+ Discipline Meetings – Used as breakouts this month to prepare for the Coordination Meeting. Documents are on SharePoint. Please share accomplishments and goals. Please have them completed by January 20.
• Geophysics: Tomorrow at 1:00 pm.
• Gas: Tomorrow at 1:00 pm
• Remote Sensing: Today at 1:00 pm
• Communications and workplace culture: Wednesday at 2:00 pm. There is an outreach spreadsheet on the SharePoint form. Please enter all your outreach activities from the past year.
• Geology: Thursday at 9:30 am.
+ USGS – Matt Haney
• Thank you to everyone for getting EPAPs completed.
• There has been a temporary freeze in new volunteer and emeritus positions. Established volunteers and emeriti may continue.
• Career ladder promotions are moving ahead.
• There was an email form Jackie Verna last week regarding a new procedure for travel statements. You now submit one pdf. There is an instructional video. Be sure to sign last.
• Charles Olin sent out an email last week regarding the Entra and Intune migration. The deadline is March 1st. You will need to re-image your computer to enroll. Charles will be enrolling the shared computers soon. You will need to log into those every 30 days to avoid errors.
• Hannah Dietterich will be on annual leave this Wednesday through Friday. She will be visiting CalVO on Thursday.
• Jordan Lubbers will be on annual leave tomorrow through next Tuesday. Matt Loewen will be coving his duty sci on Monday.
• Scott Pekalib and Jackie Verna will be visiting Anchorage Feb. 17.
• Jake Lowenstern will be visiting Anchorage Feb. 18 – 20.
• Charles Olin contacted the grounds crew to have the Grace Hall north fire exit cleared of snow.
This message is posted in the AVO Logs at https://www.avo.alaska.edu/admin/logs/posts/show/454633